# FedSlate: A Federated Deep Reinforcement Learning Recommender System*

### v1

@@ I Introduction
In numerous contexts, the correlation between user behaviors yields substantial benefits. Within the financial sector, for instance, customers often engage with both stock trading and online payment services. A collaborative model trained by these services can effectively ascertain a customer’s risk profile, investment patterns, and spending behaviors, thereby facilitating tailored financial product recommendations that meet individual needs. Similarly, in the realm of online advertising, user interactions with multiple platforms reveal interconnected behaviors. A user may explore health foods on a social network while simultaneously shopping for wellness products on an e-commerce site. Utilizing FL, platforms can collectively develop a model that captures the user’s interests and buying inclinations, enhancing the precision of targeted advertising. Consequently, a recommendation system underpinned by FL presents an approach that optimizes privacy, minimizes communication overhead, and improves long-term value for users.
@@
    We introduce FL to address the issue of interdependence in user behavior due to the influence of recommendations from different platforms in content recommendation systems. By harnessing user behavior data for collective training from multiple sources, we enhance the performance of existing reinforcement learning-based recommendation algorithms. Our approach strategically balances privacy concerns with reduced communication overhead.
@@
    Prior research on FL in recommendation systems mainly focused on applying this technique to conventional algorithms, as evidenced by [13, 14, 15, 16]. These methods typically average local model outputs without accounting for user heterogeneity [17], which can be problematic given the diverse and imbalanced nature of user preferences. Such systems are often inadequate in reflecting long-term user interests. To overcome these limitations, we introduce FedSlate, a novel algorithm that marries the concepts of SlateQ and FL. FedSlate is adept at tracking long-term user behavior and assimilating the impact of cross-platform recommendations on user activity within a singular platform. By leveraging both local and global models and adopting a “Q-value sharing” technique, FedSlate facilitates the central server in generating comprehensive Q-values for optimal content recommendation.
@@
    We propose an innovative update mechanism specific to FL augmented with reinforcement learning in our FedSlate algorithm. Unlike existing adaptive personalization strategies in FL, FedSlate utilizes a unique sequential updating process. The global network is updated multiple times based on the global Q-network’s outputs, while the local networks receive a singular update following the local agent’s environment rewards. This approach significantly enhances the algorithm’s efficiency and output.
  4. 4.
    The proposed FedSlate framework is skilled at deriving recommendation strategies from data inaccessible to local agents, an attribute of significant relevance in practical implementations. This aspect is crucial in contexts where privacy issues or data governance policies restrict information access.
@@ II Related Work
FL involves training statistical models directly on devices to develop a joint model capable of generating data across distributed nodes [10]. Traditional distributed learning methods generally presuppose that local data samples are independently and identically distributed (IID); however, FL typically operates under the premise that data among clients is non-IID [39, 40]. Prior research has primarily utilized RL to enhance FL’s performance [41]. For instance, [41] employs a Deep Q-Learning (DQL) strategy to select devices for participation in successive communication rounds, thus minimizing the number of required rounds. Related work can be found in [42, 43]. In contrast, [44] applies a deep reinforcement learning (DRL) technique to modulate the CPU frequency of faster devices within an FL training cohort, balancing energy efficiency with training velocity. Distinct from these studies, our FedSlate algorithm “federates” DRL rather than simply applying RL to optimize FL. Concerning federated recommender systems, current studies [13, 14, 15, 16], have mainly adapted FL for traditional recommendation frameworks, setting them apart from our DRL-centric FedSlate approach. The most significant influence on our work is Federated deep Reinforcement Learning (FedRL) [12], which introduces an innovative DRL paradigm to collaboratively construct high-quality models for agents. Our approach adopts the “Q-value sharing” concept from FedRL to monitor long-term user behaviors and evaluate the impact of recommended content from various platforms on a user’s actions within a specific platform, thus tackling the complexities of applying RL recommendation algorithms in a multi-platform context. While FedRL advocates for a decentralized scheme, our FedSlate algorithm is designed with centralization in mind but can also be adapted to a decentralized format for practical applications.
@@ III-A An Extended MDP Model for Slate Recommendation
We devised a user case where an individual user navigates two distinct platforms to demonstrate our algorithm, a scenario that is scalable to more platforms. However, the MDP model as initially conceived does not accommodate a user engaging with multiple platforms. To address this limitation, we propose an augmented model involving platforms A and B, wherein the MDP framework is applied independently to each. Concretely, this involves discrete recommendation agents for each platform (agent $\alpha$ for platform A and agent $\beta$ for platform B), with corresponding user states ($S_{\alpha}$ and $S_{\beta}$), agent actions ($A_{\alpha}$ and $A_{\beta}$), transition probabilities ($P_{\alpha}(s^{\prime}|s,A)$ and $P_{\beta}(s^{\prime}|s,A)$), rewards ($R_{\alpha}$ and $R_{\beta}$), and Q-functions ($Q_{\alpha}$ and $Q_{\beta}$).
@@ III-B Necessary Assumptions
  • A4: After consuming item $i$, the user’s interest transfers to other platforms.
  • A5: Only the output values of $Q_{\alpha}$ and $Q_{\beta}$ are shared for learning the joint policy $\pi^{\ast}_{fed}$. Other information, including $D_{\alpha}$ and $D_{\beta}$, is locally visible only.
A1 is the original assumption in SlateQ (SC). A2 extends the RTDS assumption in SlateQ to the multi-platform scenario. A3, A4, and A5 are specific assumptions for the single-user multi-platform context. A3 is the most important premise of our method, which is intuitive and easily acceptable. Note that A4 is a simplification of actual user behavior since regular users often switch platforms after consuming a “series” of items. However, we believe this simplification brings about concise algorithmic descriptions, and we consider A4 to be easily relaxed (by fixing the states of platforms where the user is not present and allowing state transitions to continue on the user’s current platform). A5 ensures that information about the user on different platforms is not leaked (we believe that even for the same user, information should not be shared between platforms without user authorization).
@@ III-C Recommendation Problem
After extending the original MDP model, we can formally define our recommendation problem based on A1-A5. The existing platforms $\alpha$ and $\beta$ take turns randomly recommending slates to customers and recording transitions. This results in a series of transitions $D_{\alpha}=\left\{\left\langle s_{\alpha},A_{\alpha},s{}^{\prime}_{\alpha},r_{\alpha}\right\rangle\right\}$ for agent $\alpha$ and transitions $D_{\beta}=\left\{\left\langle s_{\beta},A_{\beta}\right\rangle\right\}$ for agent $\beta$, where $D_{\alpha}$ and $D_{\beta}$ are one-to-one correspondence. Our goal is to learn a joint policy $\pi^{\ast}$ that, based on $s_{\alpha}$ and $s_{\beta}$, maximizes the lifetime value (LTV) across all platforms. It should be noted that platform B does not record the user’s response, i.e., $D_{\beta}$ does not contain $r_{\beta}$, which deviates from the previous MDP model. We adopt this setting because certain platforms may not have direct access to user feedback on recommended content (although the platform’s impact on the user is real). User preferences or aversions may be reflected on other platforms. We aim to demonstrate the friendliness of FedSlate towards these “unavailable feedback” platforms—even if a platform cannot directly update its recommendation strategy based on user feedback, it can still benefit from performance improvements in the federation.
@@ IV-A SlateQ Algorithm
In Section 3.1, we presented the MDP model for the recommendation problem. In the case of a single user and a single platform MDP, the original SlateQ algorithm aims to find an optimal policy $\pi^{\ast}$ that satisfies $\pi^{\ast}\left(s\right)=argmax_{A\in\mathcal{A}}Q^{\ast}\left(s,A\right)$. However, in Slate recommendation problems, the action space becomes extremely large, resulting in excessive computational and time resource requirements. Specifically, if we attempt to select k items from a set of $\mathcal{I}$ items to form a slate and consider the impact of the position of recommended items within the slate on user feedback, the action space would be of size ${A_{\mathcal{I}}^{k}}$.
@@
Please note that SlateQ assumes the user choice model $P(i|s,A)$ is known. Models such as MNL, CL, and cascade can easily be learned using user response data, and this does not depend on LTV.
@@ IV-B The FedSlate Algorithm
Where $|B|$ represents the batch size during training, $Y_{\alpha}=r_{\alpha}+\gamma\mathop{max}\limits_{A_{\alpha}^{\prime}\in\mathcal{A_{\alpha}}}\sum_{j\in A_{\alpha}^{\prime}}P(j|s_{\alpha}^{\prime},A_{\alpha}^{\prime})Q_{\alpha}^{f}(s_{\alpha}^{\prime},j)$ denotes the target Q-value. It is crucial to note that in Eq.(15), the updates of $Q_{\alpha}$ and $Q_{\beta}$ are both contingent upon $r_{\alpha}$ due to agent $\beta$ lacking access to $r_{\beta}$.
@@
In addition, as mentioned by [45], in real-world multi-agent systems, team rewards suffer from sparsity, making it difficult for algorithms to learn a successful team strategy to enhance overall reward. In our setting, if the rewards from Platform A and Platform B are too sparse (in other words, the information received by users on Platform A does not affect their reactions on Platform B), it is not possible to train multiple agents simultaneously using the reward from Platform A. Therefore, our algorithm has a simple variant to address this issue: Algorithm 1 is duplicated to replace Algorithm 2, and some minor modifications are made to Algorithm 3. Please refer to Algorithm 4 for the specific changes made to Algorithm 3.
By simple extension, our algorithm is capable of handling the sparsity issue of team rewards. However, it should be noted that the extended algorithm requires platform B to have access to rewards as well.
@@ V-B Algorithm Evaluation
Metric Method Environment
N=10,n=3 N=100,n=10 N=500,n=10
ETROR SlateQ N/A 3400 2020
FedSlate N/A 2340 1700
Optimal Reward SlateQ N/A 1115.072 1117.974
FedSlate N/A 1106.57 1107.092
Mean Reward Rand 947.087 942.75 938.665
@@
Environment
Metric Method N=10,n=3 N=100,n=10 N=500,n=10
Optimal Reward FedSlate 1115.281 1102.269 1129.96
Mean Reward Rand 948.294 944.982 939.979
@@
Figure 8: Comparison of Performance between Extended FedSlate Algorithm and Original Algorithm.
@@
Environment(N=100,n=10)
Metric Method dense sparse
FedSlate-Alpha 2340 3540
FedSlate(exp)-Alpha N/A 2280
FedSlate-Beta 2800 2940
ETROR FedSlate(exp)-Beta N/A 780
FedSlate-Alpha 1115.072 1081.422
FedSlate(exp)-Alpha N/A 1091.515
FedSlate-Beta 1102.269 1042.926
Optimal Reward FedSlate(exp)-Beta N/A 1108.855
@@ V-D Evaluation of Extended Algorithms
The constructed scenario requires the collaborative efforts of multiple agents to enhance the collective lifetime value of individual users across different platforms. However, in the implementation process, we used the rewards from Platform A to reflect the overall lifetime value. If the team rewards are too sparse, it may adversely affect the performance of FedSlate. To investigate the impact of this factor, we simulated the scenario of sparse team rewards by setting different random seeds for the sub-environments within the main environment, and conducted experiments under the settings of $N=100,n=10$. Additionally, we included the extended version of FedSlate for comparative analysis. The experimental results are presented in Table IV. We observed that when the team rewards in the environment are sparse, the learning process for agents in FedSlate to acquire recommendation policies takes longer, and the resulting policies are less effective. The extended algorithm of FedSlate outperforms the original version in terms of ETROR and Optimal Reward, and we observed that this improvement is more pronounced for agent $\beta$.The visual results of the comparative experiments are shown in Fig. 8. It should be noted that the extended algorithm of FedSlate is merely a supplementary approach for the original algorithm under specific circumstances. It requires agent $\beta$ to have access to reward information, which contradicts the fundamental principles of FedSlate based on vertical FL.

### v2

@@ I Introduction
In numerous contexts, the correlation between user behaviors yields substantial benefits. Within the financial sector, for instance, customers often engage with both stock trading and online payment services. A collaborative model trained by these services can effectively ascertain a customer’s risk profile, investment patterns, and spending behaviors, thereby facilitating tailored financial product recommendations that meet individual needs. Similarly, in the realm of online advertising, user interactions with multiple platforms reveal interconnected behaviors. A user may explore health foods on a social network while simultaneously shopping for wellness products on an e-commerce site. Utilizing FL, platforms can collectively develop a model that captures the user’s interests and buying inclinations, enhancing the precision of targeted advertising. Consequently, a FL-based recommender system offers a potential approach that optimizes user privacy and enhances long-term value (LTV). However, previous research on FL in recommendation systems has primarily focused on conventional algorithms with simple averaging of local model outputs [11, 12, 13, 14]. These approaches often fail to account for user heterogeneity [15], which is particularly problematic given the inherently diverse and imbalanced nature of user preferences, potentially resulting in systems that inadequately capture long-term user interests.
@@
    We propose FedSlate, a novel algorithm that integrates SlateQ methodology with federated learning principles, enabling effective monitoring of long-term user behavior patterns while incorporating cross-platform recommendation effects.
@@
    We design an innovative update mechanism that implements sequential updating protocol, involving multiple updates of the global network while maintaining single updates for local networks, thereby enhancing both algorithmic efficiency and output quality.
@@
    We demonstrate FedSlate’s effectiveness in developing recommendation strategies using data not directly accessible to local agents, addressing practical challenges in privacy-restricted scenarios.
@@ II-A Recommender Systems
II-B Federated Learning
FL involves training statistical models directly on devices to develop a joint model capable of generating data across distributed nodes [10]. Traditional distributed learning methods generally presuppose that local data samples are independently and identically distributed (IID); however, FL typically operates under the premise that data among clients is non-IID [39, 40]. Prior research has primarily utilized RL to enhance FL’s performance [41]. For instance, [41] employs a Deep Q-Learning (DQL) strategy to select devices for participation in successive communication rounds, thus minimizing the number of required rounds. Related work can be found in [42, 43]. In contrast, [44] applies a deep reinforcement learning (DRL) technique to modulate the CPU frequency of faster devices within an FL training cohort, balancing energy efficiency with training velocity. Distinct from these studies, our FedSlate algorithm “federates” DRL rather than simply applying RL to optimize FL. Concerning federated recommender systems, current studies [11, 12, 13, 14], have mainly adapted FL for traditional recommendation frameworks, setting them apart from our DRL-centric FedSlate approach. The most significant influence on our work is Federated deep Reinforcement Learning (FedRL) [17], which introduces an innovative DRL paradigm to collaboratively construct high-quality models for agents. Our approach adopts the “Q-value sharing” concept from FedRL to monitor long-term user behaviors and evaluate the impact of recommended content from various platforms on a user’s actions within a specific platform, thus tackling the complexities of applying RL recommendation algorithms in a multi-platform context. While FedRL advocates for a decentralized scheme, our FedSlate algorithm is designed with centralization in mind but can also be adapted to a decentralized format for practical applications. Please note that although our framework involves a central party, this central party does not have full transparency in accessing information from various platforms. The information that the central party can obtain is strictly limited to the Q-values generated by each platform, rather than sensitive user privacy data. Similarly, the data transmitted by the central party to each party is confined to the processed Q-values. In the FL process, such information exchange is inevitable, and there is essentially no difference between transmitting it to discrete parties and transmitting it to a central party, especially when communication is with a trusted central entity that can ensure data protection. We introduced a centralized architecture to better construct an algorithmic abstraction, thereby formally aligning with FL’s pioneering FedAvg algorithm [10]. In reality, the central party is an abstract rather than an instantiated concept, meaning that any platform instance can serve as the central party.
@@ III-A An Extended MDP Model for Slate Recommendation
To better reflect real-world usage patterns, we consider scenarios where users interact with multiple platforms simultaneously. Our algorithm is demonstrated in this multi-platform context.
However, the standard MDP framework proves insufficient for modeling cross-platform user behavior. To address this limitation, we propose an enhanced MDP framework encompassing platforms A and B, where individual MDPs operate independently on each platform. Concretely, this involves discrete recommendation agents for each platform (agent $\alpha$ for platform A and agent $\beta$ for platform B), with corresponding user states ($S_{\alpha}$ and $S_{\beta}$), agent actions ($A_{\alpha}$ and $A_{\beta}$), transition probabilities ($P_{\alpha}(s^{\prime}|s,A)$ and $P_{\beta}(s^{\prime}|s,A)$), rewards ($R_{\alpha}$ and $R_{\beta}$), and Q-functions ($Q_{\alpha}$ and $Q_{\beta}$).
@@ III-B Necessary Assumptions
  • A4: Only the output values of $Q_{\alpha}$ and $Q_{\beta}$ are shared for learning the joint policy $\pi^{\ast}_{fed}$. Other information, including transitions $D_{\alpha}=\left\{\left\langle s_{\alpha},A_{\alpha},s{}^{\prime}_{\alpha},r_{\alpha}\right\rangle\right\}$ and $D_{\beta}=\left\{\left\langle s_{\beta},A_{\beta}\right\rangle\right\}$, is locally visible only.
A1 is the original assumption in SlateQ (Single choice (SC)). A2 extends the Reward/transition dependence on selection (RTDS) assumption in SlateQ to the multi-platform scenario. A3 and A4 are specific assumptions for the single-user multi-platform context. A3 is the most important premise of our method, which is intuitive and easily acceptable. A4 ensures that information about the user on different platforms is not leaked (we believe that even for the same user, information should not be shared between platforms without user authorization).
@@ III-C Recommendation Problem
After extending the original MDP model, we can formally define our recommendation problem based on A1-A4. The existing platforms A and B take turns randomly recommending slates to customers and recording transitions. This results in a series of transitions $D_{\alpha}$ for agent $\alpha$ and transitions $D_{\beta}$ for agent $\beta$, where $D_{\alpha}$ and $D_{\beta}$ are one-to-one correspondence. Our goal is to learn a joint policy $\pi^{\ast}$ that, based on $s_{\alpha}$ and $s_{\beta}$, maximizes the lifetime value (LTV) across all platforms. It should be noted that platform B does not record the user’s response, i.e., $D_{\beta}$ does not contain $r_{\beta}$, which deviates from the previous MDP model. We adopt this setting because certain platforms may not have direct access to user feedback on recommended content (although the platform’s impact on the user is real). User preferences or aversions may be reflected on other platforms. We aim to demonstrate the friendliness of FedSlate towards these “unavailable feedback” platforms—even if a platform cannot directly update its recommendation strategy based on user feedback, it can still benefit from performance improvements in the federation.
@@ IV-A SlateQ Algorithm
In Section III-A, we presented the MDP model for the recommendation problem. For a single-user, single-platform MDP, the original SlateQ algorithm aims to find an optimal policy $\pi^{\ast}$ that satisfies $\pi^{\ast}\left(s\right)=argmax_{A\in\mathcal{A}}Q^{\ast}\left(s,A\right)$. However, in slate recommendation problems, the action space grows exponentially, imposing substantial computational and temporal complexity. Specifically, selecting and ordering $k$ items from a set of $\mathcal{I}$ items into an ordered slate results in an action space of cardinality $P(\mathcal{I},k)=\frac{|\mathcal{I}|!}{(|\mathcal{I}|-k)!}$, where $P(\mathcal{I},k)$ denotes the number of $k$-permutations of $\mathcal{I}$.
@@
Please note that SlateQ assumes the user choice model $P(i|s,A)$ is known. Models such as multinomial logit model (MNL) [45], conditional logit model (CL), and cascade model [46, 47] can easily be learned using user response data, and this does not depend on LTV.
@@ IV-B The FedSlate Algorithm
Where $|B|$ represents the batch size during training, $Y_{\alpha}=r_{\alpha}+\gamma\mathop{max}\limits_{A_{\alpha}^{\prime}\in\mathcal{A_{\alpha}}}\sum_{j\in A_{\alpha}^{\prime}}P(j|s_{\alpha}^{\prime},A_{\alpha}^{\prime})Q_{\alpha}^{f}(s_{\alpha}^{\prime},j)$ denotes the target Q-value. In Eq.(14) and Eq.(15), the updates of $Q_{\alpha}$ and $Q_{\beta}$ both depend on $r_{\alpha}$ since agent $\beta$ has no access to $r_{\beta}$.
@@
In addition, as mentioned by [48], in real-world multi-agent systems, team rewards suffer from sparsity, making it difficult for algorithms to learn a successful team strategy to enhance overall reward. In our setting, if the rewards from Platform A and Platform B are too sparse (in other words, the information received by users on Platform A does not affect their reactions on Platform B), it is not possible to train multiple agents simultaneously using the reward from Platform A. Therefore, our algorithm has a simple variant to address this issue: Algorithm 1 is duplicated to replace Algorithm 2, and some minor modifications are made to Algorithm 3. Please refer to Algorithm 4 for the specific changes made to Algorithm 3. By simple extension, our algorithm is capable of handling the sparsity issue of team rewards. However, it should be noted that the extended algorithm requires platform B to have access to rewards as well.
Although our method’s description focuses on a single user navigating content across multiple platforms, RL-based recommendation algorithms are indeed applicable in industrial-level systems. Such systems generally comprise three key components [1]: a user feedback log system; offline DNN training using these logs (typically for scoring user-item pairs); and online recommendation services that perform real-time scoring and ranking to select the top-k items. FedSlate, a Q-learning based algorithm optimized for long-term benefits, constructs its state space with features that predict immediate user feedback rewards and are capable of self-prediction. These characteristics allow for the reuse of existing log systems. Addressing user generalization, FedSlate adapts the MDP for all users, leveraging the generalization capabilities already present in existing systems. Regarding user feedback modeling, FedSlate utilizes existing ranking system definitions, such as predicted Click-Through Rate (pCTR), allowing the system to employ these conventional models to predict the likelihood of an item being selected.
@@ V-B Algorithm Evaluation
Metric Method Environment N=10,n=3 N=100,n=10 N=500,n=10 ETROR SlateQ[1] N/A 3400 2020 FedSlate N/A 2340 1700 Optimal Reward SlateQ N/A 1115.072 1117.974 FedSlate N/A 1106.57 1107.092 Mean Reward Rand 947.087 942.75 938.665
@@
Environment Metric Method N=10,n=3 N=100,n=10 N=500,n=10 Optimal Reward FedSlate 1115.281 1102.269 1129.96 Mean Reward Rand 948.294 944.982 939.979
@@
Figure 8: Comparison of Performance between Extended FedSlate Algorithm and Baseline Algorithm.
@@
Metric Method Environment(N=100,n=10) dense sparse ETROR FedSlate-Alpha 2340 3540 FedSlate(exp)-Alpha N/A 2280 FedSlate-Beta 2800 2940 FedSlate(exp)-Beta N/A 780 Optimal Reward FedSlate-Alpha 1115.072 1081.422 FedSlate(exp)-Alpha N/A 1091.515 FedSlate-Beta 1102.269 1042.926 FedSlate(exp)-Beta N/A 1108.855 Mean Reward FedQ[49] 981.971 973.150
@@ V-D Evaluation of Extended Algorithms
The proposed scenario requires collaborative efforts among multiple agents to optimize the collective LTV of individual users across different platforms. In our implementation, we used Platform A’s rewards as a proxy for overall lifetime value. However, sparse team rewards can potentially degrade FedSlate’s performance. To investigate this effect, we simulated sparse team reward scenarios by implementing different random seeds for sub-environments within the main environment, conducting experiments with parameters $N=100$ and $n=10$.
Our experimental framework encompasses both the original FedSlate algorithm and an extended version designed to address sparse team rewards. While a comprehensive comparison with state-of-the-art methods would ideally position our approach within the current technological landscape, the integration of FL into recommendation systems remains an emerging field. Our literature review [11, 51, 52] identified the FedQ [49] algorithm, which leverages global recommendation system insights, as the most relevant comparable work. However, FedQ primarily focuses on applying FL to traditional recommendation models without addressing cross-platform LTV considerations. Due to FedQ’s requirement for complete user feedback, we constructed a controlled environment with universal access to user feedback for comparative analysis with our extended FedSlate version.
Table IV presents the experimental results. In environments with sparse team rewards, FedSlate agents required extended learning periods to develop recommendation policies, with diminished policy effectiveness. The extended FedSlate algorithm demonstrated superior performance compared to the original version in both ETROR metrics and Optimal Reward calculations, with agent $\beta$ showing particularly marked improvements. Fig. 8 illustrates the comparative experimental outcomes. A significant limitation of the FedQ algorithm lies in its inability to account for cross-platform LTV, leading it to optimize for immediate rewards during training. This approach often results in recommending short-term engaging but potentially harmful items (metaphorically termed ‘chocolate’ items), ultimately reducing long-term user engagement and necessitating policy relearning. In the “Choc vs. Kale” scenario, FedQ’s performance barely exceeded random recommendation benchmarks and exhibited significant instability, underscoring FedSlate’s distinctive advantages among current methodologies.
V-E Extended Experiments Across Multiple Platforms
(a) Configuration 1
(b) Configuration 2
Figure 9: Performance evaluation of FedSlate algorithm across expanded platform configurations. The designation ‘FedSlate-1’ represents Platform 1’s implementation within the federated learning framework. Configuration 1 (5 platforms with user feedback access) exhibited enhanced convergence and robustness in recommendation policy learning relative to baseline metrics. Configuration 2 demonstrated that platforms without direct feedback access (Platforms 4 and 5) achieved performance parity with the original SlateQ algorithm through federated knowledge transfer from Platforms 1-3, establishing the framework’s scalability and sustained effectiveness with increased federation size.
To comprehensively evaluate the proposed FedSlate framework, we conducted extended experiments under two distinct configurations. In the first configuration, five platforms with complete access to their respective user feedback participated in the study. This setup employed the extended FedSlate algorithm described in Section IV-B to investigate potential additional benefits derived from cross-platform feedback sharing. The second configuration involved five platforms, with two platforms lacking access to user feedback, enabling us to assess the framework’s scalability and effectiveness with increased participation.
The experimental parameters, including hyperparameters, remained consistent with those detailed in Section V-B, with results presented in Figure 9. Analysis of the first configuration revealed that all participating platforms demonstrated accelerated and more robust learning of recommendation strategies. Furthermore, leveraging cross-platform user LTV metrics resulted in consistently superior recommendation performance compared to baseline methods, as evidenced by increased reward metrics.
In the second configuration, platforms 4 and 5, which lacked direct access to user feedback, achieved performance levels comparable to the original SlateQ algorithm through indirect benefits from platforms 1, 2, and 3. This outcome demonstrates the framework’s sustained effectiveness with increased participant numbers. While Section V-B noted that FedSlate might impact the optimal local strategies of feedback-enabled agents despite improving their training efficiency, the results from the second configuration suggest this limitation diminishes with increased federation participation. This improvement can be attributed to the synergistic effects of collaborative user interest development, where the fundamental components of user engagement yield greater benefits as the number of participating platforms increases, thereby enhancing the overall effectiveness of content recommendations.